OpenStudy (anonymous):

can some one plz help i will give a medal and a fan!!! b-15=-12

OpenStudy (texaschic101):

add 15 to both sides and you will find b

OpenStudy (anonymous):

hold up a min plz

OpenStudy (texaschic101):

b - 15 = - 12 + 15 + 15 -------------- b = -12 + 15......can you finish ?

OpenStudy (anonymous):